A company with working capital of $214,679 and a current ratio of 1.8 pays a $75,097 short-term liability. Determine the amount of working capital immediately after payment. Select the correct answer. $289,776 $214,679 $139,582 $75,097A company with working capital of $214,679 and a current ratio of 1.8 pays a $75,097 short-term liability. Determine the amount of working capital immediately after payment. Select the correct answer. $289,776 $214,679 $139,582 $75,097A company with working capital of $214,679 and a current ratio of 1.8 pays a $75,097 short-term liability. Determine the amount of working capital immediately after payment. Select the correct answer. $289,776 $214,679 $139,582 $75,097A company with working capital of $214,679 and a current ratio of 1.8 pays a $75,097 short-term liability.
Answer #1
Working Capital = $214,679
Current Ratio = 1.8
Working Capital = Current Assets-Current Liabilities
Let Working Capital = WC, Current Assets= CA, Current Liabilitires = CL
Current Ratio= Current Assets/ Current Liabilties
Current Ratio= CA/CL
1.8= CA/CL
CA= 1.8CL
WC = CA-CL
      214,679.00 = CA-CL
      214,679.00 = 1.8CL- CL ( From above CSA= 1.8 CL)
      214,679.00 = .8 CL
CL = 214679/.8
CL                                  268,348.75
CA = 268348.75*1.8 (CA= 1.8CL)
CA                                  483,027.75
After payment of short term liability of $ 75097, the current Liability will be
CL = 268348.75-75097
193251.8
Working Capital 483027.75-193251.8
Working Capital $289,776
The working capital will be $ 289776